OpenStudy (anonymous):

Solve the Equations Beloww ; 2/5n+6=10 -7-6y=13

OpenStudy (anonymous):

2/5n+6=10 -6 -6 2/5n=16 ---- --- 2/5 2/5 n=40

OpenStudy (anonymous):

-7-6y=13 +7 +7 -6y=20 --- -- -6 -6 y= -20/6